A soccer ball is kicked from the ground with an initial speed of 19.2 m/s at an upward angle of 44.7˚. A player 50.1 m away in the direction of the kick starts running to meet the ball at that instant. What must be his average speed if he is to meet the ball just before it hits the ground? Neglect air resistance.Use g=9.81 m/s2


The tolerance is +-1 in the 3rd significant digit,


1
Expert's answer
2021-11-18T15:09:02-0500
"t=2\frac{19.2\sin{44.7}}{9.81}=2.574\ s\\v=\frac{50.1}{2.574}=19.5\frac{m}{s}"